Fifteen years from now Darren will be four times as old as he was three years ago. How old is Darren now?
Answer by BurningFlame_marc(77)

Hello!
I will show you how I solve these kinds of problems.


Representation:
Let x-3 be the age of darren 3 years ago, x+15 be his age 15 years from now.
Equation:
x+15=4(x-3)
Solution:
x+15=4(x-3)
x=4x-12-15
x=4x-27
-3x=-27
3x=27
x=9

Final Answer:
He is now 9 years old...